OpenStudy (anonymous):

***WILL MEDAL AND FAN*** A 10" diameter pumpkin pie is cut into six equal servings. What is the area of the top of each piece of pie? a) 3.33pi inches b) 4.17pi inches c) 16.66pi inches

OpenStudy (anonymous):

im thinking its A cause im trying to work it out but it just comes close to ten

OpenStudy (phi):

10 inch diameter means 5 inch radius Do you know how to find the area of a circle with r=5 ?

OpenStudy (anonymous):

Yea. The area would be 78.5 right?

OpenStudy (anonymous):

@phi

OpenStudy (phi):

I would not make it decimal (because the answer choices have "pi" in them can you find the area, but do not change pi to a decimal?

OpenStudy (anonymous):

Hmmm`

OpenStudy (phi):

first, what is the formula for Area of a circle?

OpenStudy (anonymous):

pi x radius squared

OpenStudy (phi):

and r= 5 r squared ( \( r^2 \) ) = 5*5= 25 so Area= 25 * 3.14159265358979323846264338327950288419716939937510582097494459230781640628620899862803482534211706798214808651328230664709384460955058... pi goes on forever... rather than truncating to approximately 3.14 (which is close but not exactly correct, as you can see), and rather than typing all those numbers (which you can't because they go on forever), people just say... call that number pi Area= \(25 \pi\) and let's not multiply it out unless you want a decimal number... and you can use as many digits of pi as makes sense for your problem (i.e. how accurate you want to be)

OpenStudy (phi):

so the area of the circle is 25pi now divide that area into 6 equal pieces. what is the area of 1/6 of the whole circle? can you do that ? (and leave pi alone)

OpenStudy (anonymous):

It's 4.6

OpenStudy (phi):

it's \[ \frac{25\pi}{6} \] which is the same as \[ \frac{25}{6} \pi \] (hopefully that makes sense... if not you should learn why you can do that ) now what is 25/6 ?

OpenStudy (anonymous):

4.16

OpenStudy (phi):

you mean 4.16666... (the 6 repeats) if you round to the nearest 100th (2nd decimal to the right of the decimal point) you get 4.17 now you can answer the question

OpenStudy (anonymous):

Oh thanks so much. and btw, can you answer another question for me? Just one more/

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Please

OpenStudy (phi):

Please make a new post.

OpenStudy (anonymous):

I will